# CLQCD Collaboration

### v2

@@ Acknowledgments
Then the optimal operator which can suppress the current-involved ESC can be expressed as a linear combination ${\cal N}^{\rm opt}={\cal N}+c^{\rm opt}{\cal N}_{X}$ with the coefficient $c$ to be determined. In our calculation, we use the interpolation operator ${\cal N}_{X}$ directly and not to project the reducible lattice representations onto the spin-$1/2$, isospin-$1/2$ subspace to obtain irreducible operators Prelovsek:2016iyo; Detmold:2024ifm. With our choice ${\cal N}_{X}$, excitations with higher spin or isospin are allowed and are handled effectively by varying the coefficient $c$ in the linear combination. In particular, both ${\cal N}_{S}$ and ${\cal N}_{T}$ also create isospin-$3/2$ states, which share the same energy with isospin-$1/2$ states because we take $m_{u}=m_{d}$. Moreover, ${\cal N}_{T}$ couples to both spin-$1/2$ and spin-$3/2$ $N\,b_{1}$ states; these correspond to different linear combinations of polarization components but are degenerate eigenstates of the Hamiltonian. Consequently, such higher-spin or higher-isospin excited states share the same Euclidean-time dependence in the linear error correction and can be absorbed by a rescaling of $c$.
@@
Figure 7: The dependence of the source-sink separation $t_{f}$ for the ratios $\mathcal{R}_{X}^{\rm mid}(t_{f},c)\equiv\mathcal{R}_{X}(t_{f},t=t_{f}/2;{\cal N}(c))$ and Feynman-Hellmann fnunctions $\mathcal{R}_{X}^{\text{FH}}(t_{f})$ for the tensor charge $X=T$ on various ensembles: C48P23 (upper left), C32P23 (upper right), C48P14 (lower left), F48P21 (lower right), which cover different pion masses, volumes, and lattice spacings. The gray band represents the result from the joint three-state fit. The minimum source-sink separation for fitting is $t_{f}^{\text{min}}\sim 0.3~\text{fm}$.
@@
On each ensemble, we can find an optimized coefficient $c=c^{\rm opt}$ with mild time-dependencies in $R^{\rm mid}_{X}(t_{f})$ and $R^{\rm FH}_{X}(t_{f})$ in moderate and large $t_{f}$. Fig. 7 displays the results for the tensor charge $g_{T}$ on four characteristic ensembles covering different pion masses (including the physical pion mass), volumes, and lattice spacings:
@@
3) C48P14 with a=0.105 fm, $m_{\pi}\sim 135$ MeV, $L\sim$ 5.1 fm (lower left panel);
4) F48P21 with a=0.077 fm, $m_{\pi}\sim 210$ MeV, $L\sim$ 3.7 fm (lower right panel);
@@
We also test the influence from the nucleon-pion scattering state Barca:2022uhi; Grebe:2023tfx; Alexandrou:2024tin using a two-operator GEVP basis, $\{{\cal N},{\cal N}\pi\}$. The optimized linear combination is obtained by solving Eq. (13), where the stabilized solutions can be obtained at relatively large $t^{\prime}$ and $t_{0}$. As tested on C48P14 and shown in Fig. 11, the results with and without GEVP have similar time dependencies in moderate and large $t_{f}$, indicating that the nucleon-pion scattering state has a smaller influence on the ESC for both $g_{T}$ and $g_{S}$, compared to the current-involved excited states. Therefore, ${\cal N}\pi$ is not considered in extracting the ground state matrix elements in our calculations.
@@
Previous studies employing similar five-quark interpolating operators (e.g., Alexandrou:2024tps; Wang:2023omf; Barca:2024hrl; Barca:2025det) included only the first three contributions, omitting the $\mathcal{N}_{X}\rightarrow\mathcal{N}_{X}$ term due to its exceptionally complex contraction structure. As summarized in Table 2, the number of required quark diagrams increases substantially: from just 2 for the two-point function of the standard three-quark interpolation operator $\cal N$}, to 3 times more for its three-point function with $O_{X}=\bar{u}\Gamma_{X}u-\bar{d}\Gamma_{X}d$, and similarly for the $\mathcal{N}_{X}\rightarrow\mathcal{N}$ two-point function. The $\mathcal{N}_{X}\rightarrow\mathcal{N}_{X}$ two-point function requires 4.3 times more diagrams than the standard case, with an additional 4.3-fold increase for its three-point function, making the computation of the $\mathcal{N}_{X}\rightarrow\mathcal{N}_{X}$ contribution extremely challenging even at the propagator contraction stage.
@@
$N_{\rm e}$ $N_{\rm st}$ $g_{V}^{\rm 2-state}$ $g_{S}^{\rm 2-state}$ $g_{S}^{\rm 3-state}$ $g_{T}^{\rm 2-state}$ $g_{T}^{\rm 3-state}$
C24P34 50 250 0.9927(62) 1.026(18) 1.025(23) 1.0055(57) 1.0040(66)
C24P29 50 50 0.9943(26) 1.022(27) 1.017(37) 0.9838(36) 0.9810(45)
C32P29 60 60 0.9961(23) 1.113(13) 1.077(13) 0.9971(20) 0.9879(55)
C24P23 50 50 0.9903(35) 0.880(24) 0.942(33) 0.9661(40) 0.9670(39)
C32P23 60 60 0.9978(38) 1.052(29) 1.033(19) 0.9816(37) 0.9739(38)
C48P23 100 200 0.9950(17) 1.108(19) 1.100(18) 0.9949(36) 0.9822(58)
C48P14 150 200 0.9961(55) 1.027(30) 1.048(34) 0.9756(69) 0.9723(81)
C64P14 150 50 0.9967(24) 1.097(25) 1.108(27) 0.9846(85) 0.9865(75)
E32P29 50 50 0.9980(52) 1.037(30) 1.004(28) 1.0035(45) 1.0017(63)
F32P30 100 200 0.9964(10) 0.943(24) 0.997(40) 0.9990(28) 0.9994(37)
F48P30 100 200 0.9987(40) 1.117(15) 1.084(17) 1.0210(24) 1.0151(35)
F32P21 50 50 0.9992(57) 0.897(34) 0.906(42) 0.9643(55) 0.9644(38)
F48P21 100 60 0.9974(23) 1.005(19) 1.045(26) 0.9910(26) 0.9855(33)
F64P13 140 60 1.0010(26) 0.975(20) 0.988(30) 0.9853(79) 1.0009(54)
G36P29 60 140 0.9995(38) 0.967(32) 0.951(35) 1.0088(60) 1.0066(61)
H48P32 100 200 0.9981(09) 0.963(28) 0.977(36) 1.0201(47) 1.0217(53)
$g_{V/S/T}^{\rm physical}$ 0.9998(12) 1.094(27) 1.107(32) 1.0234(54) 1.0256(58)
.3 Systematic uncertainties
@@
For the continuum, chiral and infinite-volume extrapolations, we use the Akaike information criterion (AIC) to estimate the systematic error from different extrapolation models, with the AIC weight we use here are Borsanyi:2020mff:
@@
.3.1 Continuum extrapolation
For continuum extrapolation, we include the following models in the AIC:
@@
Figure 14: The lattice spacing dependence of the $g_{S}$ and $g_{T}$ with data points corrected to physical pion mass and infinite volume limits. The data points at the same lattice spacing but different pion mass and volume are horizontally shifted to improve readability.
Fig. 14 presents the continuum extrapolation for $g_{T}$ (blue dots and band) and $g_{S}$ (orange dots and band), using the default $a^{2}$ ansatz. The data shown have been corrected for finite-volume effects and unphysical pion masses using the parameters determined from the global fit. As shown in the figure, results at the five lattice spacings display a clear linear $a^{2}$ dependence with no evident $a^{4}$ corrections. The slopes for $g_{S}$ and $g_{T}$, however, exhibit opposite signs.
.3.2 Chiral extrapolation
@@
.3.3 Infinite volume extrapolation

### v3

@@ Acknowledgments
Then the optimal operator which can suppress the current-involved ESC can be expressed as a linear combination ${\cal N}^{\rm opt}={\cal N}+c^{\rm opt}{\cal N}_{X}$ with the coefficient $c$ to be determined. In our calculation, we use the interpolation operator ${\cal N}_{X}$ directly and not to project the reducible lattice representations onto the spin-$1/2$, isospin-$1/2$ subspace to obtain irreducible operators Prelovsek et al. (2017); Detmold et al. (2024). With our choice ${\cal N}_{X}$, excitations with higher spin or isospin are allowed and are handled effectively by varying the coefficient $c$ in the linear combination. In particular, both ${\cal N}_{S}$ and ${\cal N}_{T}$ also create isospin-$3/2$ states, which share the same energy with isospin-$1/2$ states because we take $m_{u}=m_{d}$. Moreover, ${\cal N}_{T}$ couples to both spin-$1/2$ and spin-$3/2$ $N\,b_{1}$ states; these correspond to different linear combinations of polarization components but are degenerate eigenstates of the Hamiltonian. Consequently, such higher-spin or higher-isospin excited states share the same Euclidean-time dependence in the linear error correction and can be absorbed by a rescaling of $c$.
@@
Figure 7: The dependence of the source-sink separation $t_{f}$ for the ratios $\mathcal{R}_{X}^{\rm mid}(t_{f},c)\equiv\mathcal{R}_{X}(t_{f},t=t_{f}/2;{\cal N}(c))$ and Feynman-Hellmann fnunctions $\mathcal{R}_{X}^{\text{FH}}(t_{f})$ for the tensor charge $X=T$ on various ensembles: C48P23 (upper left), C32P23 (upper right), C48P14 (middle left), F48P21 (middle right),F48P30 (lower left), F48P30 (lower right), which cover different pion masses, volumes, and lattice spacings. The gray band represents the result from the joint three-state fit. The minimum source-sink separation for fitting is $t_{f}^{\text{min}}\sim 0.3~\text{fm}$.
@@
On each ensemble, we can find an optimized coefficient $c=c^{\rm opt}$ with mild time-dependencies in $R^{\rm mid}_{X}(t_{f})$ and $R^{\rm FH}_{X}(t_{f})$ in moderate and large $t_{f}$. Fig. 7 displays the results for the tensor charge $g_{T}$ on six characteristic ensembles covering different pion masses (including the physical pion mass), volumes, and lattice spacings:
@@
3) C48P14 with a=0.105 fm, $m_{\pi}\sim 135$ MeV, $L\sim$ 5.1 fm (middle left panel);
4) F48P21 with a=0.077 fm, $m_{\pi}\sim 210$ MeV, $L\sim$ 3.7 fm (middle right panel);
5) F32P30 with a=0.077 fm, $m_{\pi}\sim 300$ MeV, $L\sim$ 2.4 fm (lower left panel);
6) F48P30 with a=0.077 fm, $m_{\pi}\sim 300$ MeV, $L\sim$ 3.7 fm (lower right panel);
@@
Figure 11: Effective masses on the C48P14 ensemble with physical pion mass, before (left panel) and after (right panel) solving the GEVP in Eq. (13) with the three-operator basis $\{{\cal N},({\cal N}\pi)_{\rm P-wave},({\cal N}\pi\pi)_{\rm S-wave}\}$. The grey, blue and red band shows the energy of nucleon ground state, non-interacting nucleon-pion threshold and nucleon-pion-pion threshold, respectively.
To assess the potential influence of nucleon-pion scattering states, we perform a GEVP analysis using a three-operator basis $\{{\cal N},({\cal N}\pi)_{\rm P-wave},({\cal N}\pi\pi)_{\rm S-wave}\}$ on the physical-pion-mass ensemble C48P14. Figure 11 compares the effective masses before (left panel) and after (right panel) solving the GEVP. The gray, blue, and red bands indicate the energies of the nucleon ground state, the non-interacting $N\pi$ threshold, and the $N\pi\pi$ threshold, respectively. After applying the GEVP, the resulting eigenstates align well with these thresholds, confirming that the method successfully isolates the expected low-lying states.
For the influence on $g_{S,T}$, we just using a two-operator GEVP basis, $\{{\cal N},({\cal N}\pi)_{\rm P-wave}\}$ since that involving ${\cal N}\pi\pi)_{\rm S-wave}$ is extremely complicated. The optimized linear combination is obtained by solving Eq. (13), where the stabilized solutions can be obtained at relatively large $t^{\prime}$ and $t_{0}$. As tested on C48P14 and shown in Fig. 12, the results with and without GEVP have similar time dependencies in moderate and large $t_{f}$, indicating that the nucleon-pion scattering state has a smaller influence on the ESC for both $g_{T}$ and $g_{S}$, compared to the current-involved excited states. Therefore, ${\cal N}\pi$ is not considered in extracting the ground state matrix elements in our calculations.
@@
Previous studies employing similar five-quark interpolating operators (e.g., Alexandrou et al. (2025); Wang et al. (2024); Barca et al. (2025); Barca (2025)) included only the first three contributions, omitting the $\mathcal{N}_{X}\rightarrow\mathcal{N}_{X}$ term due to its exceptionally complex contraction structure. As summarized in Table 3, the number of required quark diagrams increases substantially: from just 2 for the two-point function of the standard three-quark interpolation operator $\cal N$}, to 3 times more for its three-point function with $O_{X}=\bar{u}\Gamma_{X}u-\bar{d}\Gamma_{X}d$, and similarly for the $\mathcal{N}_{X}\rightarrow\mathcal{N}$ two-point function. The $\mathcal{N}_{X}\rightarrow\mathcal{N}_{X}$ two-point function requires 4.3 times more diagrams than the standard case, with an additional 4.3-fold increase for its three-point function, making the computation of the $\mathcal{N}_{X}\rightarrow\mathcal{N}_{X}$ contribution extremely challenging even at the propagator contraction stage.
@@
aa(fm) $n_{L}^{3}\times n_{T}$ $m_{\pi}$(MeV) $m_{\pi}L$ $m_{\eta_{s}}(\rm MeV)$ $n_{\rm cfg}$ interval $\tau$ $N_{\rm e}$ $N_{\rm st}$ $g_{V}^{\rm 2-state}$ $g_{S}^{\rm 2-state}$ $g_{S}^{\rm 3-state}$ $g_{T}^{\rm 2-state}$ $g_{T}^{\rm 3-state}$
[…]
$g_{V/S/T}^{\rm physical}$ 0.9998(12) 1.098(27) 1.106(31) 1.0236(52) 1.0264(53)
In this section, we present a comprehensive overview of our simulation setup as well as the methodology adopted for autocorrelation analysis. The Hybrid Monte Carlo (HMC) algorithm is employed for generating gauge field configurations. The integration step size $\tau$ and the interval between configurations for each ensemble are summarized in Table 4. To further enhance the efficiency of our simulations, we implement the Hasenbusch mass preconditioning technique Hasenbusch (2001), which effectively accelerates the HMC algorithm by improving the conditioning of the fermion matrix and thus increasing the acceptance rate.
The thermal equilibrium of our Markov chain are carefully checked. In Fig. 14, we display the evolution of the plaquette value $\langle P\rangle=u_{0}^{4}$ and topological charge along the HMC trajectories for the F48P21, C48P14, and H48P32 ensembles. As illustrated, both the plaquette and the topological charge fluctuate randomly around their respective mean values, demonstrating the absence of long-term drifts. This behavior provides a clear indication that the generated configurations are well equilibrated and that our ensembles have reached thermal equilibrium.
A notable advantage of the blending method lies in its ability to extract significantly more information from a single configuration by approximating the all-to-all propagator. Consequently, far fewer configurations are required compared to traditional techniques. However, when using fewer configurations, it becomes essential to properly account for auto-correlations present along the Markov chain trajectory. To quantitatively evaluate the impact of auto-correlation in this context, we perform a thorough statistical analysis, focusing on the variance of binned means as outlined in Ref. Chang et al. (2018).
Building upon the auto-correlation studies of basic observables such as $u_{0}$, $Q$, $m_{\pi}$, and $m_{\eta_{c}}$ in Ref. Du et al. (2025), we conduct a dedicated binning analysis on our data. Specifically, we compute $R_{S,T}^{\rm mid}(t_{f})$ using different bin sizes $n=\{1,2,3,4\}$ for both the $c=0$ and $c=c_{\text{opt}}$ interpolators across three representative ensembles:
  • F48P21: $a=0.077\,\mathrm{fm}$, $m_{\pi}\approx 210\,\mathrm{MeV}$, $m_{\pi}L=3.91$,
  • F64P13: $a=0.077\,\mathrm{fm}$, $m_{\pi}\approx 130\,\mathrm{MeV}$, $m_{\pi}L=3.37$,
  • H48P32: $a=0.052\,\mathrm{fm}$, $m_{\pi}\approx 320\,\mathrm{MeV}$, $m_{\pi}L=4.00$.
For each case, the mean value and standard deviation are estimated using a sufficiently large number of bootstrap resamples. In the absence of significant autocorrelations, the standard deviation should remain nearly constant as the bin size $n$ varies. The dependence of both the mean value and standard deviation on bin size $n$ is depicted in Fig. 15. The results show that the standard deviation remains stable for all tested ensembles, indicating that autocorrelation effects are negligible in our data. As a result, the statistical errors reported in this study can be considered robust and unaffected by autocorrelation.
Figure 14: The evolution of the plaquette and topological charge along the HMC trajectory. present the plaquette(upon three figures) and topological charge(down three figures) as the function of the configuration number.
Figure 15: The autocorrelation analysis of $g_{\rm S,T}$ is shown by illustrating $R_{S,T}^{\rm min}(t_{f})$ with $c=0$ and $c=c_{\text{opt}}$ with bin size is $\{1,2,3,4\}$(from left to right, color from dark to shallow) on the F48P21 (left panels), F64P13 (middle panels) and H48P32 (right panels) ensembles, for the tensor (upper panels) and scalar (lower panels) cases. The crosses represent the results for $R^{\rm min}$ with $c=0$, while the circles represent the results with $c=c_{\rm opt}$.
.4 The sensitivity analysis of fit window
In this Appendix, we present the fit windows used to extract $g_{\rm S,T}$ on each of the three ensembles, along with the corresponding $\chi^{2}/\mathrm{d.o.f.}$ values in Table 5. To systematically assess the stability of our results, we performed a dedicated sensitivity analysis by varying the start ($t_{\rm min}$) and end ($t_{\rm max}$) of the fit window by $\pm$1 lattice spacing relative to the values listed in Table 5. This test was conducted on three representative ensembles, F48P21, F64P13 and H48P32, with the results compiled in Table 6.
Our analysis demonstrates the robustness of the extraction: on all the three ensembles, the fitted values of $g_{S}$ and $g_{T}$ remain consistent across all tested window variations. As expected, statistical uncertainties moderately increase (decrease) when $t_{\rm min}$ is increased ($t_{\rm max}$ is decreased), confirming that our chosen window represents a sound balance between statistical precision and systematic control.
Therefore, our chosen fit windows (documented in Table 5) were specifically selected to satisfy two key criteria: (1) ensure the stability and good constraint of the three-state fit parameters, and (2) achieve an optimal balance between statistical precision and systematic control over excited-states. This rigorous sensitivity analysis confirms that our final results are robust against reasonable variations in the fit window.
Figure 16: The ratio of the original data to the fit results for $R_{S}(t,t_{f})$ and $R_{T}(t,t_{f})$ as a function of $t-t_{f}/2$ on the F64P13 ensemble. The hollow and solid data points correspond to $c=0$ and $c=c_{\rm opt}$, respectively.
Fig. 16 displays the normalized fit residuals—the original data with statistical uncertainty divided by the central value of the fit function, for both the two-point correlators and the three-point ratios $R_{S}(t,t_{f})$ and $R_{T}(t,t_{f})$ on the F64P13 ensemble, plotted as a function of the time separation.
The ratios are consistently scattered around unity across the entire fitted range and remain well within the uncertainty. This demonstrates visually that our multi-state fit model provides an excellent, unbiased description of the correlator data, with no discernible systematic trends.
Table 5: The fitting information for $g_{S}$ and $g_{T}$ from the two-state and three-state fits.
Ensemble $n_{\text{state}}$ $g_{S}$ $g_{T}$
$t_{\text{min}}(a)$ $t_{\text{max}}(a)$ $g_{S}$ $\chi^{2}$/d.o.f $t_{\text{min}}(a)$ $t_{\text{max}}(a)$ $g_{T}$ $\chi^{2}$/d.o.f
C24P34 2 2 8 1.036(18) 0.57 6 10 1.0182(58) 0.02
3 2 8 1.038(23) 0.04 3 12 1.0167(67) 0.04
C24P29 2 3 10 1.027(29) 0.63 6 10 0.9933(37) 0.10
3 3 12 1.027(37) 0.06 3 12 0.9904(45) 0.23
C32P29 2 3 10 1.126(13) 0.98 6 10 1.0099(20) 0.17
3 3 10 1.091(13) 0.19 3 12 1.0007(55) 0.34
C24P23 2 2 8 0.880(24) 0.84 6 10 0.9661(40) 0.09
3 3 8 0.942(33) 0.03 3 10 0.9670(39) 0.09
C32P23 2 6 12 1.058(29) 0.09 6 10 0.9873(37) 0.19
3 3 12 1.039(19) 0.07 3 12 0.9796(38) 0.28
C48P23 2 3 12 1.115(19) 1.30 6 10 1.0008(36) 0.33
3 3 12 1.107(19) 0.01 3 12 0.9881(59) 0.27
C48P14 2 3 8 1.029(30) 0.58 6 10 0.9777(69) 0.16
3 3 8 1.050(34) 0.02 3 12 0.9743(81) 0.33
C64P14 2 4 10 1.099(25) 1.52 6 10 0.9866(85) 0.29
3 4 10 1.111(27) 0.15 6 12 0.9885(75) 0.66
E32P29 2 4 14 1.046(30) 0.70 7 12 1.0121(45) 0.06
3 4 14 1.012(29) 0.11 4 14 1.0103(64) 0.14
F32P30 2 6 16 0.954(26) 1.21 8 14 1.0054(28) 0.30
3 6 16 1.004(40) 0.12 4 17 1.0059(37) 0.01
F48P30 2 6 16 1.124(16) 0.69 8 14 1.0276(24) 0.45
3 6 16 1.091(17) 0.06 4 13 1.0217(35) 0.31
F32P21 2 3 11 0.904(34) 1.26 8 14 0.9683(56) 0.16
3 3 11 0.910(42) 0.12 4 13 0.9685(38) 0.21
F48P21 2 4 16 1.084(26) 0.55 8 14 0.9951(26) 0.42
3 4 16 1.049(26) 0.04 4 13 0.9897(33) 0.28
F64P13 2 6 11 0.990(23) 0.42 10 17 0.9927(64) 0.93
3 3 12 0.997(29) 0.16 3 15 0.9982(39) 0.83
G36P29 2 5 12 0.972(32) 0.43 9 16 1.0142(61) 0.08
3 5 12 0.956(35) 0.01 5 19 1.0120(61) 0.16
H48P32 2 6 18 0.919(19) 0.55 12 21 1.0243(47) 0.11
3 6 18 0.982(36) 0.03 6 23 1.0260(53) 0.43
Table 6: Analysis of the sensitivity of $g_{S}$ and $g_{T}$ to the choice of fit window.
Ensemble $n_{\text{state}}$ $g_{S}$ $g_{T}$
$t_{\text{min}}(a)$ $t_{\text{max}}(a)$ $g_{S}$ $\chi^{2}/d.o.f$ $t_{\text{min}}(a)$ $t_{\text{max}}(a)$ $g_{T}$ $\chi^{2}/d.o.f$
F48P21 3 4 16 1.049(26) 0.04 4 13 0.9897(33) 0.28
3 3 16 1.048(24) 0.05 3 13 0.9900(33) 0.43
3 5 16 1.050(26) 0.04 5 13 0.9890(35) 0.23
3 4 15 1.051(23) 0.02 4 12 0.9913(31) 0.25
3 4 17 1.047(27) 0.09 4 14 0.9876(39) 0.31
F64P13 3 3 12 0.997(29) 0.16 3 15 0.9982(39) 0.83
3 2 12 1.000(29) 0.18 2 15 1.0003(47) 1.15
3 4 12 0.992(32) 0.15 4 15 0.9974(41) 0.79
3 3 11 1.004(26) 0.12 3 14 0.9990(37) 0.82
3 3 13 0.993(31) 0.16 3 16 0.9968(42) 0.91
H48P32 3 6 18 0.982(36) 0.03 6 23 1.0260(53) 0.43
3 5 18 0.982(42) 0.03 5 23 1.0258(48) 0.46
3 7 18 0.981(36) 0.03 7 23 1.0261(69) 0.41
3 6 17 0.986(33) 0.02 6 22 1.0258(50) 0.43
3 6 19 0.973(38) 0.05 6 24 1.0261(57) 0.42
.5 Systematic uncertainties
@@
For the continuum, chiral and infinite-volume extrapolations, we use the Akaike information criterion (AIC) to estimate the systematic error from different extrapolation models, with the AIC weight we use here are Borsanyi et al. (2021):
@@
.5.1 Continuum extrapolation
For clover fermions, a general $O(a)$-improved operator takes the form $O^{\rm imp}=(1+c_{0}m_{q}a)O+c_{1}aO^{\rm D}$, where $O^{\rm D}$ is a higher-dimensional operator containing a derivative. Perturbative calculations Capitani et al. (2001) show that, after accounting for cancellations between quark-field and operator corrections, $c_{0}=1+\mathcal{O}(\alpha_{s})$ and $c_{1}=\mathcal{O}(\alpha_{s})$. In our analysis, the $O(a)$ effects (specifically the $O(m_{q}a)$ terms) for the scalar and tensor currents are removed by the renormalization procedure. This is achieved by combining the vector charge renormalization constant $Z_{V}$ (determined from a hadron matrix element at the unitary light quark mass) with the ratios $Z_{S/T}/Z_{V}$ taken in the chiral limit.
Consequently, the leading residual discretization errors are of order $O(a\alpha_{s})$. As demonstrated even for the charm quark case in Ref. Du et al. (2025), these effects are numerically small. Thus we include the following models in the AIC for continuum extrapolation:
@@
Figure 18: The lattice spacing dependence of the $g_{S}$ and $g_{T}$ with data points corrected to physical pion mass and infinite volume limits using the default $a^{2}$ ansatz. The data points at the same lattice spacing but different pion mass and volume are horizontally shifted to improve readability.
Fig. 18 presents the continuum extrapolation for $g_{T}$ (blue dots and band) and $g_{S}$ (orange dots and band), using the default $a^{2}$ ansatz. The data shown have been corrected for finite-volume effects and unphysical pion masses using the parameters determined from the global fit. As shown in the figure, results at the five lattice spacings display a clear linear $a^{2}$ dependence with no evident $a^{4}$ corrections. The slopes for $g_{S}$ and $g_{T}$, however, exhibit opposite signs. Fig. 19 displays the continuum extrapolation using six additional models, with each subfigure containing two fit lines corresponding to two distinct models. As shown in the figures, higher-order terms ($a^{3}$ or $a^{4}$) do not significantly affect the extrapolated value. Including the $a\alpha_{s}$ term or restricting the data to finer lattice spacings increases the uncertainty considerably, while the central value remains largely unchanged.
Figure 19: The lattice spacing dependence of the $g_{S}$ and $g_{T}$ with data points corrected to physical pion mass and infinite volume limits. The data points at the same lattice spacing but different pion mass and volume are horizontally shifted to improve readability. For the three figures, the left panel presents the extrapolation models $a^{2}+a^{3}$ and $a\log(u_{0})+a^{2}+a^{3}$; the middle panel shows the extrapolation models $a^{2}+a^{3}+a^{4}$ and $a\log(u_{0})+a^{2}+a^{3}+a^{4}$; and the right panel displays the extrapolation model $a^{2}$ using only ensembles with $a<0.1$ fm, along with $c_{a}^{(1)}a\log(u_{0})+c_{a}^{(2)}a^{2}$. The darker fit line corresponds to the model that includes the $a\log(u_{0})$ term.
.5.2 Chiral extrapolation
@@
.5.3 Infinite volume extrapolation
@@
Figure 22: Comparison of this work versus other collaborations at finite volume, with all data extrapolated to the continuum limit.
Fig. 22 further plots $g_{T}$ and $g_{S}$, corrected for discretization and chiral extrapolation effects, against $e^{-m_{\pi}L}$. It shows that our high-precision data (blue triangles) are fully consistent within $2\sigma$ with the results from RQCD Bali et al. (2023) (orange circles) and PNDME Jang et al. (2024) (green crosses) at comparable volumes. The statistical uncertainties of the earlier results are substantially larger, which prevented those studies from definitively identifying the FV trend or discriminating between different model ansatzes.